# Fostering Intrinsic Motivation in Reinforcement Learning with Pretrained Foundation Models

### v1

@@ 1 Introduction
Inspired by human psychology, intrinsic motivation encourages agents to be curious (i.e., seek novel situations) in the environment, fostering a more robust and adaptive learning process. This typically involves generating an intrinsic reward $r^{i}_{t}$ that quantifies the novelty or surprise, which is then added to the actual extrinsic reward from the environment $r^{e}_{t}$. This addition leads to an updated reward signal $r_{t}=r^{e}_{t}+\beta\cdot r^{i}_{t}$, wherein $\beta\in\mathbb{R}^{+}$ balances between exploiting known information in the environment with fostering the agent’s curiosity. However, the effectiveness of these methods often depends on the quality of the state representations used to compute the novelty bonus, which are commonly learned during the agent’s training process.
@@ 4 Experimental Setup
Experiments are conducted over the MiniGrid benchmark, a collection of procedurally generated grid-based environments that challenge RL agents to explore and navigate effectively in the presence of reward sparsity. In MiniGrid, the environment information can come in two different forms: a compact, encoded vector that summarizes key information about the grid ; and an RGB image that provides a detailed visual representation. Specifically, we evaluate performance across MultiRoom (MN7S4, MN7S8, MN12S10), KeyCorridor (KS3R3, KS4R3), and ObstructedMaze (O2Dlh) environments.
@@
MultiRoom-N7-S4 [10M] 0.9M 2.71M 0.57M 0.48M
MultiRoom-N7-S8 [20M] - - - 2.65M
MultiRoom-N12-S10 [30M] - - - 3.76M
KeyCorridorS3R3 [30M] 4.01M 4.19M 2.27M 4.95M
KeyCorridorS4R3 [50M] - 16.19M 8.59M 10.84M
ObstructedMaze-2Dlh [50M] 23.34M 19.53M 17.92M 20.36M
@@ 5 Results
Table 1 provides the number of steps required for RIDE and FoMoRL to converge to the optimal policy in different MiniGrid environments, considering both partial observations and full state information. For both approaches, the use of full state information significantly augments the sample-efficiency. In complex MultiRoom environments only FoMoRL with full state access manages to effectively train. However, in KeyCorridor and ObstructuredMaze, RIDE performs slightly better than FoMoRL. We hypothesize that this is because these environments contain a higher diversity of objects, allowing RIDE to effectively distinguish between different states and learn a better representation. In contrast, FoMoRL relies on pretrained embeddings from CLIP, which remain frozen during training and may lack the adaptability required to fully capture the variations in these more diverse environments.
@@
Table 2: Performance of RIDE and FoMoRL in terms of the number of required steps to achieve convergence, without considering the episodic term. That is, $\forall s_{t}$ $\xrightarrow{}$ $\sqrt{N_{ep}(s_{t})}=1$ in Eqs (1),(2).
@@
MultiRoom-N7-S4 [20M] 3.07M 14.24M
MultiRoom-N7-S8 [30M] - 9.61M
MultiRoom-N12-S10 [50M] - 20.15M
KeyCorridorS3R3 [50M] 4.15M 5.76M
KeyCorridorS4R3 [50M] 11.04M -
ObstructedMaze-2Dlh [80M] 22.52M 19.64M
@@ Appendix A Model Architectures and Hyperparameters
Table 4: $\beta$ coefficient without episodic counts.
@@
We used a discount factor $\gamma=0.99$, a clipping factor $\epsilon=0.2$, 4 epochs per training step, and $\lambda=0.95$ for GAE. We employed 16 parallel environments to gather rollouts of size 128, resulting in a total horizon of 2,048 steps between updates.

### v2

@@ 1 Introduction
Inspired by human psychology, intrinsic motivation encourages agents to be curious (i.e., seek novel situations) in the environment, fostering a more robust and adaptive learning process Aubret et al. (2019). This typically involves generating an intrinsic reward $r^{i}_{t}$ that quantifies the novelty or surprise, which is then added to the actual extrinsic reward from the environment $r^{e}_{t}$. This addition leads to an updated reward signal $r_{t}=r^{e}_{t}+\beta\cdot r^{i}_{t}$, wherein $\beta\in\mathbb{R}^{+}$ balances between exploiting known information in the environment with fostering the agent’s curiosity. However, the effectiveness of these methods often depends on the quality of the state representations used to compute the novelty bonus, which are commonly learned during the agent’s training process.
@@ 4 Experimental Setup
Experiments are conducted over the MiniGrid benchmark, a collection of procedurally generated grid-based environments that challenge RL agents to explore and navigate effectively in the presence of reward sparsity. In MiniGrid, the agent can receive environment information in one of two ways: as a full state, where the agent has complete access to the entire grid layout, $s_{t}$, or as a partial observation, $o_{t}$, where the agent’s view is restricted to a limited, egocentric field centered on its current position. Partial observations are thus a subset of the full state, unable to capture the areas behind walls or closed doors. The information, whether full state or partial observation, can be represented in two forms: as a compact encoded vector (e.g., ${s_{t}^{\text{enc}},o_{t}^{\text{enc}}}$) that summarizes the environment’s key elements, or as an RGB image (e.g., ${s_{t}^{\text{rgb}},o_{t}^{\text{rgb}}}$).
Figure 1: Illustration of full state and partial observation formats in the MiniGrid environment, with two types of representations: encoded and RGB. The top row shows the full state $s_{t}^{enc}$, where the agent has access to the entire grid layout, represented as a three-channel encoded matrix. The three channels encode object type, color, and state (e.g., open or closed doors, agent orientation). The dotted pink square in this row indicates the region that would compose the agent’s observation in a partial view. The middle row displays this partial observation $o_{t}^{enc}$, where the agent’s perception is limited to a $7\times 7$ egocentric field centered on its position. The green square in this row highlights the area visible to the agent, showing how walls and other objects limit the agent’s field of vision. Lastly, the bottom row shows the RGB representation of a partial observation $o_{t}^{rgb}$, where each cell is represented according to three color channels (red, green, and blue) for a pixel-based view.
Figure 1 illustrates these different observation types, showing examples of both encoded and RGB representations, as well as full and partial observations. We evaluate performance across several MiniGrid environments, including MultiRoom (MN7S4, MN7S8, MN12S10), KeyCorridor (KS3R3, KS4R3), and ObstructedMaze (O2Dlh).
@@
MultiRoom-N7-S4 [10M] 1.53M 4.03M 0.93M 0.58M
MultiRoom-N7-S8 [20M] - - - 2.8M
MultiRoom-N12-S10 [30M] - - - 4.06M
KeyCorridorS3R3 [30M] 6.82M 5.29M 2.07M 5.13M
KeyCorridorS4R3 [50M] - 17.06M 7.92M 8.15M
ObstructedMaze-2Dlh [50M] 39.03M 19.92M 29.02M 16.53M
@@ 5 Results
Table 1 provides the number of steps required for RIDE and FoMoRL to converge to the optimal policy in different MiniGrid environments, considering both partial observations and full state information. For both approaches, the use of full state information significantly augments the sample-efficiency. In complex MultiRoom environments only FoMoRL with full state access manages to effectively train. However, in KeyCorridor, RIDE performs slightly better than FoMoRL. We hypothesize that this is because these environments contain a higher diversity of objects, allowing RIDE to effectively distinguish between different states and learn a better representation. In contrast, FoMoRL relies on pretrained embeddings from CLIP, which remain frozen during training and may lack the adaptability required to fully capture the variations in these more diverse environments.
@@
Table 2: Performance of RIDE and FoMoRL in terms of the number of required steps to achieve convergence, without considering the episodic term. That is, $\forall o_{t}$ $\xrightarrow{}$ $\sqrt{N_{ep}(o_{t})}=1$ in Expressions (1) and (2).
@@
MultiRoom-N7-S4 [20M] 3.19M 14.92M
MultiRoom-N7-S8 [30M] - 11.89M
MultiRoom-N12-S10 [50M] - 25.32M
KeyCorridorS3R3 [50M] 5.27M 7.12M
KeyCorridorS4R3 [50M] 11.95M -
ObstructedMaze-2Dlh [80M] 24.76M 15.54M
We refer to Appendix B for more details about the results summarized in Tables 1 and 2.
@@ Appendix A Model Architectures and Hyperparameters
Table 4: $\beta$ coefficient without episodic counts, i.e., $\forall o_{t}$ $\xrightarrow{}$ $\sqrt{N_{ep}(o_{t})}=1$ in Equations (1) and (2).
@@
We used a discount factor $\gamma=0.99$, a clipping factor $\epsilon=0.2$, 4 epochs per training step, and $\lambda=0.95$ for GAE. We employed 16 parallel environments to collect rolloutsS of size 128, resulting in a total horizon of 2,048 steps between updates.
@@
Appendix B Graphical Results
Figure 2: Performance comparison between RIDE (orange) and FoMoRL (blue) across various MiniGrid environments and input types. Each subfigure presents the average return over training steps, illustrating learning progress and convergence speed for each algorithm in specific settings. Solid lines represent full observations ($s_{t}$), while dotted lines denote partial observations ($o_{t}$). The horizontal black dashed line indicates the expected return of the optimal policy. The shaded area represents the standard deviation computed across 3 different seeds.